# Phenomenology of Non-Abelian Gauge and Goldstone Bosons in a U(2) Flavor Model

### v1

@@ 6 Summary and outlook
In this study, we have explored the implications of the $U(2)_{F}=SU(2)_{F}\times U(1)_{F}$ flavor model proposed in [32], with a focus on the phenomenology of the bosons associated with $SU(2)_{F}$. Both the global and local $SU(2)_{F}$ scenarios entail the existence of new particles — PNGBs or gauge bosons, respectively — with distinct interactions and phenomenological properties that imply stringent constraints on the flavor symmetry breaking scale.
@@
Our results demonstrate that the existence of light NP particles (which is unavoidable in the case of global flavor symmetries but possible also in the local case) implies that laboratory experiments can test the flavor breaking scale, that is, the origin of fermion masses and mixing, up to ultra-high energies, in some cases even beyond the limits set by astrophysical observations.
As an outlook, we can mention a couple of directions that deserve further investigation. For instance, it would be interesting to assess the capability of future gravitational wave (GW) observatories to test the high-energy symmetry breaking regime of this model, similarly to what done in [22] for $U(1)_{F}$ Froggatt-Nielsen models. A possibly observable stochastic GW background (SGWB) would result from cosmic strings produced at the phase transition temperature $T\approx v_{\phi}$, which requires the reheating temperature $T_{\text{RH}}$ at the end of inflation to satisfy $T_{\text{RH}}>v_{\phi}$. This would imply that, unlike the scenario considered above, axiflavon DM is produced through misalignment after inflation and domain walls cannot be diluted by inflation. The distinctive predictions, possible issues and resulting imprints on the SGWB following from this setup deserve dedicated future studies. Another interesting direction is related to the cosmological role of the PNGBs $\pi^{\prime}$ or the gauge bosons $W^{\prime}$ if they are ultra-light — a natural situation in the former case, perhaps less in the latter. Even if additional DM candidates seem not to be required within this model, such ultra-light fields — if substantially produced in the early universe — could act as dark radiation or be constrained by big-bang nucleosynthesis inducing further interesting signatures or tests of the model.
@@ Appendix B Experimental limits
As discussed in Section 5.1, for light weakly-coupled bosons, the main constraints arise from searches for an invisible particle in flavour-violating meson and lepton decays. The corresponding limits reported in Table 2 show the case of a practically massless boson $m_{X}\approx 0$. In addition, the experimental limits on $\mu\to e\,X_{\text{inv}}$ are sensitive to the chirality of the coupling as muon experiments employ polarized beams, as discussed in details in [35]. The displayed limit refers to a light boson with $m_{X}\approx 0$ and substantial couplings with RH lepton currents (like in the case under study). In order to take into account the $m_{X}$ dependence for heavier bosons, we employ the average limit $\text{BR}(\mu\to e\,X_{\text{inv}})<6\times 10^{-6}$ on bosons with RH interactions obtained in [62]. For the other processes, we adopt the exclusion curves reported in [75] for $\tau\to\ell\,X_{\text{inv}}$ (see also [82]), that on $K^{+}\to\pi^{+}\,X_{\text{inv}}$ from [17] (that also includes the limit on $\pi^{0}\to\text{inv}$ [83] in the mass gap $m_{X}\approx m_{\pi^{0}}$ where the former search is not sensitive due to the large $K^{+}\to\pi^{+}\pi^{0}$ background), and finally the $2\sigma$ exclusion on $B^{+}\to K^{+}\,X_{\text{inv}}$ estimated in [54] based on the Belle II measurement of $B^{+}\to K^{+}\nu\bar{\nu}$ [82].
@@ Appendix C Astrophysical limits
The most relevant astrophysical constraint on couplings to quarks also comes from neutron star cooling during supernova explosions. Translating the effective coupling to nucleons as defined in [87] to the case of our bosons $X$:
$g_{XN}\equiv\frac{2C_{N}m_{N}}{v_{\phi}}\penalty 10000\ \text{global case}\,,\quad g_{XN}\equiv\frac{1}{2}g_{F}C_{N}\quad\text{local case}\,$, (64)
@@ Appendix D Axiflavon phenomenology
where the relation between the axion decay constant and the $U(2)_{F}$ breaking scale is $f_{a}\simeq 0.17\times v_{\phi}$, see Eq. (13). which are consistent with those calculated in [32] considering that we employ a different fit to the $\mathcal{O}(1)$} coefficients and a different $U(1)_{F}$ charge assignment in the lepton sector (both from [35]). As, we can see, $K\to\pi a$ and $\mu\to\pi a$ give the most stringent constraints that, however, are orders of magnitude weaker than similar limits that we obtained for the light $SU(2)_{F}$ bosons. This is a consequence of the $U(2)_{F}$ protection of the axiflavon couplings that, in particular, suppresses flavor violation in the 1-2 sector. Therefore, more relevant are the astrophysical constraints, which read
$f_{a}\gtrsim 5\times 10^{8}\,\text{GeV}\penalty 10000\ \penalty 10000\ (\text{SN1987A})\,.\quad f_{a}\gtrsim 3.5\times 10^{8}\,\text{GeV}\penalty 10000\ \penalty 10000\ (\text{WD})\,.$ (71)

### v2

@@ 6 Summary and outlook
In this study, we have explored the implications of the $U(2)_{F}=SU(2)_{F}\times U(1)_{F}$ flavor model proposed in [42], with a focus on the phenomenology of the bosons associated with $SU(2)_{F}$. Both the global and local $SU(2)_{F}$ scenarios entail the existence of new particles — PNGBs or gauge bosons, respectively — with distinctive interactions and phenomenological properties that imply stringent constraints on the flavor symmetry breaking scale.
@@
Our results demonstrate that the existence of light new physics particles (which is unavoidable in the case of global continuous flavor symmetries but possible also in the local case) implies that laboratory experiments can test the flavor breaking scale, that is, the origin of fermion masses and mixing, up to ultra-high energies, in some cases even beyond the limits set by astrophysical observations.
As an outlook, we can mention a couple of directions that deserve further investigation. For instance, it would be interesting to assess the capability of future gravitational wave (GW) observatories to test the high-energy symmetry breaking regime of this model, similarly to what done in [32] for $U(1)_{F}$ Froggatt-Nielsen models. A possibly observable stochastic GW background (SGWB) would result from cosmic strings produced at the phase transition temperature $T\approx v_{\phi}$, which requires the reheating temperature $T_{\text{RH}}$ at the end of inflation to satisfy $T_{\text{RH}}>v_{\phi}$. This would imply that, unlike the scenario considered above, axiflavon DM is produced through misalignment after inflation and domain walls cannot be diluted by inflation. The distinctive predictions, possible issues and resulting imprints on the SGWB following from this setup deserve dedicated future studies — for studies on the SGWB from phase transitions and cosmic strings within other kinds of $SU(2)$ flavor models, see [83, 84]. Another interesting direction is related to the cosmological role of the PNGBs $\pi^{\prime}$ or the gauge bosons $W^{\prime}$ if they are ultra-light — a natural situation in the former case. Even if additional DM candidates seem not to be required within this model, such ultra-light fields — if substantially produced in the early universe — could act as dark radiation or be constrained by big-bang nucleosynthesis inducing further interesting signatures or tests of the model.
@@ Appendix B Experimental limits
As discussed in Section 5.1, for light weakly-coupled bosons, the main constraints arise from searches for an invisible particle in flavor-violating meson and lepton decays. The corresponding limits reported in Table 2 show the case of a practically massless boson $m_{X}\approx 0$. In addition, the experimental limits on $\mu\to e\,X_{\text{inv}}$ are sensitive to the chirality of the coupling as muon experiments employ polarized beams, as discussed in details in [47]. The displayed limit refers to a light boson with $m_{X}\approx 0$ and substantial couplings with RH lepton currents (like in the case under study). In order to take into account the $m_{X}$ dependence for heavier bosons, we employ the average limit $\text{BR}(\mu\to e\,X_{\text{inv}})<6\times 10^{-6}$ on bosons with RH interactions obtained in [78]. For the other processes, we adopt the exclusion curves reported in [93] for $\tau\to\ell\,X_{\text{inv}}$ (see also [100]), that on $K^{+}\to\pi^{+}\,X_{\text{inv}}$ from [26] (that also includes the limit on $\pi^{0}\to\text{inv}$ [101] in the mass gap $m_{X}\approx m_{\pi^{0}}$ where the former search is not sensitive due to the large $K^{+}\to\pi^{+}\pi^{0}$ background), and finally the $2\sigma$ exclusion on $B^{+}\to K^{+}\,X_{\text{inv}}$ estimated in [69] based on the Belle II measurement of $B^{+}\to K^{+}\nu\bar{\nu}$ [100].
For certain values of the model’s parameters, our light bosons could also decay inside the detector after traveling a macroscopic distance, hence giving rise to signatures commonly referred to as “displaced vertices”. To the best of our knowledge, searches for displaced vertices in the decays of Table 2 have never been attempted. Their viability and possible impact strongly depend on the specific process. For instance, the running Mu3e experiment is only sensitive to prompt $\mu\to ee\bar{e}$ decays, because the three electron tracks are required to reconstruct the primary vertex in order to tame accidental backgrounds, see e.g. [72]. Similarly, within $B$ factories, a secondary vertex (SV) with a significant displacement from the collision point is necessary to identify a $B$ meson decay, so that the kaon and the leptons from $B\to K\ell\ell^{\prime}$ should point to such a SV. On the other hand, for a kaon decay to be visible, it is just required that the light particle $X$ decay within the decay volume (e.g. $L\sim 100\,\mathrm{m}$ for NA62), thus it does not matter whether that occurs promptly or with some displacement. In contrast, searching for displaced $\tau$ LFV decays at Belle II is a viable possibility that might improve the sensitivity to our light particles and requires further investigation.
In general, as we have seen in our phenomenological discussion, when limits from searches for an invisible light particle $X$ are by far the strongest, the corresponding visible decays just serve to exclude the large coupling regime. Since there is no gap between the regions excluded by searches for invisible and visible $X$ to be filled, we expect that searches for displaced vertices would not lead to increased sensitivity. Again, this is not the case of $\tau$ decays, where the visible and invisible modes are currently complementary and set competing limits, as shown by our Figures 1 and 2.
@@ Appendix C Astrophysical limits
The most relevant astrophysical constraint on couplings to quarks also comes from neutron star cooling during supernova explosions. Translating the effective coupling to nucleons as defined in [105] to the case of our bosons $X$, we can write:
$g_{XN}\equiv\frac{2C_{N}m_{N}}{v_{\phi}}\penalty 10000\ \penalty 10000\ \text{(global case)}\,,\quad g_{XN}\equiv\frac{1}{2}g_{F}C_{N}\penalty 10000\ \penalty 10000\ \text{(local case)}\,$, (64)
@@ Appendix D Axiflavon phenomenology
where the relation between the axion decay constant and the $U(2)_{F}$ breaking scale is $f_{a}\simeq 0.17\times v_{\phi}$, see Eq. (13). These limits are consistent with those calculated in [42] considering that we employ a different fit to the $\mathcal{O}(1)$} coefficients and a different $U(1)_{F}$ charge assignment in the lepton sector (both from [47]). As, we can see, $K\to\pi a$ and $\mu\to\pi a$ give the most stringent constraints that, however, are orders of magnitude weaker than similar limits that we obtained for the light $SU(2)_{F}$ bosons. This is a consequence of the $U(2)_{F}$ protection of the axiflavon couplings that, in particular, suppresses flavor violation in the 1-2 sector. Therefore, more relevant are the astrophysical constraints, which read
$f_{a}\gtrsim 5\times 10^{8}\,\text{GeV}\penalty 10000\ \penalty 10000\ (\text{SN1987A})\,,\quad f_{a}\gtrsim 3.5\times 10^{8}\,\text{GeV}\penalty 10000\ \penalty 10000\ (\text{WD})\,.$ (71)